OpenStudy (anonymous):

For the given function, find the vertical and horizontal asymptote(s) (if there are any). f(x) = x-5/x^2-1

OpenStudy (agent0smith):

Vertical: find where the denominator equals zero.

OpenStudy (anonymous):

wait so its verical

OpenStudy (agent0smith):

That's how you find verticals. Horizontal: if the denominator is a higher degree than the numerator, then the asymptote is y=0

OpenStudy (anonymous):

vetical

OpenStudy (agent0smith):

...? These are how you find the vertical and horizontal asymptotes.

OpenStudy (agent0smith):

x^2-1 = 0 will find V.A. Horizontal: if the denominator is a higher degree than the numerator, then the asymptote is y=0

OpenStudy (anonymous):

x^2=1

OpenStudy (agent0smith):

x^2-1 is a difference of two squares

OpenStudy (agent0smith):

(x+1)(x-1) = 0

OpenStudy (anonymous):

x = 5, y = 0

OpenStudy (agent0smith):

you forgot (x+1)(x-1) = 0

OpenStudy (agent0smith):

x=5 is a zero or x-intercept. Not asymptote.

OpenStudy (anonymous):

x = 1, x = -1, y = 0